OpenStudy (anonymous):

The given point is on the graph of y=f(x). Find a point on the graph of y=g(x). g(x)=f(x-1)+3; (6,16)

jimthompson5910 (jim_thompson5910):

since (6,16) is on the graph of y = f(x), we know that f(6) = 16

jimthompson5910 (jim_thompson5910):

compare f(x-1) to f(6). What must x be equal to?

jimthompson5910 (jim_thompson5910):

ie, f(x-1) = f(6) ----> x = ???

OpenStudy (anonymous):

@jim_thompson5910 okay so x=5, f(5-1)+3 = 7 so how to I find the y cordinate?

OpenStudy (anonymous):

oh I got it (7,19) very very VERY confusing tho!

jimthompson5910 (jim_thompson5910):

(7,19) is close, but not correct

jimthompson5910 (jim_thompson5910):

x = 7 g(x)=f(x-1)+3 g(7)=f(7-1)+3 g(7)=f(6)+3 g(7)=16+3 g(7)=19 nvm I misread lol. So yeah, (7,19) is on g(x)

OpenStudy (anonymous):

lol! I have another similar question... I will try on my own real quick.

jimthompson5910 (jim_thompson5910):

ok post the question and what you get

OpenStudy (anonymous):

@jim_thompson5910 g(x)=1/4f(x);(-4,20) I got (-1,5) and I know its wrong because answer keys says its (-4,5)

jimthompson5910 (jim_thompson5910):

(-4,20) means f(-4) = 20

jimthompson5910 (jim_thompson5910):

plug x = -4 into g(x) g(x)=1/4f(x) g(-4)=1/4f(-4) g(-4)=1/4*(20) g(-4)=5

jimthompson5910 (jim_thompson5910):

you're only taking 1/4 of the y coordinate while you leave the x coordinate alone

OpenStudy (anonymous):

OOOH!! I see.. thank you so much!!! :) I have a test tmrw I really appreciate your help!